OpenStudy (anonymous):

What is the solution to x + 11.4 = 13.9?

OpenStudy (anonymous):

x=11.4-13.9 x=-2.5

OpenStudy (anonymous):

@Vshal3 :You are wrong,Check it again @geargiapeach :subtract 11.4 from each side x=2.5

OpenStudy (cwrw238):

x + 11.4 = 13.9 x = 13.9 - 11.4 = 2.5

OpenStudy (anonymous):

yep,saw the equation wrongly